Three boys have marbles in the ratio 19 : 5 : 3. If the boy with the least number has 9 marbles, how many marbles does the boy with the greatest number have?
(A) 57 (B) 62 (C) 48 (D) 54 (E) 70

Found 2 solutions by checkley77, Theo:
Answer by checkley77(12844)

3:9=19:x
3x=9*19
3x=171
x=171/3
x=57 ans. for the greatest number of marbles.
ANSWER A).

Answer by Theo(13342)

let A be the one with the largest number.
let B be the one with the middle number.
let C be the one with the smallest number.

let a = the number of marbles that A has.
let b = the number of marbles that B has.
let c = the number of marbles that C has.

you get:

a:b:c = 19:5:3

C has 9 marbles.

9/3 = 3 which means that C has 3 times the marbles as it's ratio number.

to keep the ratios intact, all others in the chain have to be multiplied by 3.

you get:

a:b:c = 19:5:3 = 57:15:9

this means that A has 57 marbles while B has 15 marbles while C has 9 marbles.

you can also use the ratio formula to find the correct value.

b:c = 5:3

This is the same as b/c = 5/3 since a fraction is a ratio of 2 numbers.

you have b/c = 5/3.
you know that c = 9.
the equation becomes:

b/9 = 5/3
multiply both sides of this equation by 9 to get:
b = (9*5)/3 = 45/3 = 15.

now that you know b, you can solve for a in the same manner.

a:b = 19:5

this is the same as a/b = 19/5.

you have a/b = 19/5.
you know that b is 15.
the equation becomes:

a/15 = 19/5
multiply both sides of this equation by 15 to get:
a = (15*19)/5 = 285/5 = 57

you wind up with:

a = 57
b = 15
c = 9

your question was how many marbles does the boy with the greatest number have?

your answer is selection A (57).